DIRC Operations Report
Wednesday 2008/01/23 to Wednesday 2008/01/30
Oncall: Jennifer Prendki (Joe Schwiening on backup)
Fast Monitoring: Joe Schwiening and Jennifer Prendki
Luminosities this week
PEP peak lumi: ~9.5 x 1033
(last week: ~10.0 x 1033)
PEP delivery: 3.72 fb-1
BaBar integrated lumi: 3.47 fb-1
Run 7 total delivery: 17.67 fb-1
Run 7 total BaBar recorded: 17.13 fb-1 total,
including 0 fb-1 off-peak data and including 16.51
fb-1 on the Y(3S)
PEP activities/issues this week
- The last 2 days PEP has had problems with the vacuum (vacuum
leaks) near IR2. This prevents collisions.
BaBar activities/issues this week
- Generally high background spikes this week due to the PEP vacuum
DIRC activities/issues this week
- The main problem with the DIRC last week was due to bad
LED calibrations. This started to happen on the 18th. It was
discovered Thursday. With the help of Chris O'Grady, this
problem was traced to a broken "partition master" board in the
electronics house. We have modified the calibration code in
order to try to prevent this bad calibrations from being
written in database. The bad calibrations which were written to
the database have been removed by Nicolas.
- Yesterday we saw 3 hot channels (sector #8, DFB #7) which were giving a lot of
FX errors. Today 8 channels including those appeared to have
died. These are the same 8 channels that have died 3 times in the past ...
See DIRC logbook for details
Lumi and Background this week
- Joe and Jennifer rotating oncall and backup duties.